# Antidistillation Sampling

### v2

@@ 3 Empirical Results
To demonstrate the effectiveness of antidistillation sampling in practice, we simulate real-world distillation by instantiating separate teacher, proxy student, and student models. Concretely, we use deepseek-ai/DeepSeek-R1-Distill-Qwen-7B [25] as the teacher model, Qwen/Qwen2.5-3B [26] as the proxy model, meta-llama/Llama-3.2-3B [27] as the student model. We evaluate the teacher and the student peformance on the GSM8K [8] and MATH [7] benchmarks, as strong performance on these benchmarks requires training on high-quality traces.
Baselines.
Throughout, we compare to a straightforward sampling method: temperature sampling. In this scheme, tokens are sampled from the teacher’s distribution with some temperature $\tau$; this approximates what an API endpoint might be doing in the nominal case. We present comparisons to additional baselines in §A.
Hyperparameters.
Antidistillation sampling involves two key hyperparameters: $\epsilon$, which controls the approximation power of the finite-difference computation, and $\lambda$, which controls the degree to which the sampling distribution mixes in the antidistillation penalty. Regarding the choice of $\epsilon$, we empirically verify that the finite difference approximation in eq. 11 is close to the JVP in eq. 8 on a smaller model in §C.1. In practice, we find that $\epsilon=10^{-4}$ works reasonably well for the BFloat16 models we use. With regard to $\lambda$, we sweep over several values to study the trade off involved in perturbing the sampling distribution.
3.1 Main Results
In Figure 1, we provide evidence that antidistillation sampling effectively meets the desiderata laid out in §2.1. That is, for a fixed teacher accuracy, we observe significantly worse performance for students distilled from traces sampled via antidistillation sampling relative to models distilled from traces generated via temperature scaling. When performing distillation, we use LoRA [28] with rank 128, $\alpha$ equal to 128, dropout probability set to 0, learning rate equal to 0.0005, weight decay coefficient of 0.1, and the maximum gradient norm set to 1.0. We train with a cosine learning rate schedule with a warm-up parameter of 0.1, batch size of 32, and we train for 4 epochs.
In Figure 2, we vary $\lambda$ to examine the degree of control antidistillation sampling gives model owners over the trade-off between teacher performance and distillability. We note that as we use distinct student and proxy student architectures, our results indicate that antidistillation sampling can generalize across architectures.
@@ Appendix A Additional baselines
Appendix B Example traces
More example traces can be found at https://antidistillation.com/traces.
Appendix C Hyperparameters
C.1 Verifying finite difference approximation
We empirically verify that the finite difference in eq. 11 behaves as expected by computing the relative error between the finite difference result and term produced from autograd. As shown in fig. 4, we see it well approximates the autograd computed result for appropriately chosen step size. Here we compute $\left\langle\nabla\ell\left(\theta_{P}\right),\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})\right\rangle$ and stack the different values of $x_{t+1}$ into a $V$ dimensional vector $\widehat{\Delta}$ and compare to the autograd vector $\Delta$. We compute relative error being sensitive only to the direction as
@@ C.1 Verifying finite difference approximation
Due to memory constraint, we run this numerical experiment using GPT2-Large [29]. Here we demonstrate that the finite difference can be used to estimate the derivatives in the low precision BFloat16 format. In particular, too small an $\epsilon$ leads to round-off error in the perturbation and too large $\epsilon$ leads to high truncation error in the Taylor expansion, with a sweet spot in the middle. The actual choice of $\epsilon$ depends heavily on the model size (and numerical precision), so we recommend choosing this value on the exact model in question. In our actual experiment, we pick $\epsilon$ empirically to be $10^{-4}$.

### v3

@@ 4.1 Experimental setup
Architectures. To demonstrate the effectiveness of antidistillation sampling in practice, we simulate realistic distillation by instantiating distinct teacher, proxy student, and actual student models. Specifically, we use deepseek-ai/DeepSeek-R1-Distill-Qwen-7B (DeepSeek-AI, 2025) as the teacher model, Qwen/Qwen2.5-3B(Yang et al., 2024) as the proxy model, and meta-llama/Llama-3.2-3B(Grattafiori et al., 2024) as the student model (we examine other architecture configurations in §4.2).
[…]
Calculating the downstream loss. Calculating $\nabla\ell(\theta_{P})$ requires evaluating the proxy model on a holdout set of reasoning traces. For our experiments, we use the first 70% of our train data as the training set and the remaining 30% as the holdout set. We use the teacher to generate reasoning traces on the holdout set, and then calculate $\nabla\ell(\theta_{P})$ on these reasoning traces using gradient accumulation while masking out the system and question prompt.
Baselines. Our primary baseline is temperature sampling, which approximates standard API endpoint behavior while providing a controlled way to degrade teacher performance. Importantly, temperature sampling ensures we fairly compare trade-offs against a straightforward baseline that—like our method—degrades teacher performance by modifying the sampling procedure. One other point of comparison to this baseline is that antidistillation sampling requires two forward passes on the proxy model for each forward pass on the teacher, independent of $\lambda$. Since we choose the proxy model to be approximately half the size of the teacher, this amounts to doubling the computation needed to sample outputs from the model compared to temperature sampling. In practice, one might choose a much smaller proxy model to reduce the overhead further.
Hyperparameters. Antidistillation sampling involves two key hyperparameters: $\epsilon$, which controls the approximation power of the finite-difference computation, and $\lambda$, which determines the weight of the antidistillation penalty in the sampling distribution from eq.˜11.
  • For $\epsilon$, we empirically verify that our finite difference approximation in eq.˜10 closely matches the JVP result in eq.˜7 (see §B.1). In practice, we find that $\epsilon=10^{-2}$ works well for BFloat16 models, which is close to the minimum in Figure 7.
  • For $\lambda$, we conduct a comprehensive sweep to characterize the utility-distillability trade-off that results from perturbing the sampling distribution.
We use a max generation length of $1024$ for both GSM8K and MMLU and $2048$ for MATH. For antidistillation sampling, we use a temperature of $\tau=0.6$; we found that sweeping between $\tau\in[0,1]$ does not significantly impact antidistillation performance. All of our experiments are performed on nodes with 8 NVIDIA H100 GPUs and we use the transformers package (Wolf et al., 2020), the trl toolkit (von Werra et al., 2020), and the accelerate library (Gugger et al., 2022).
Distillation protocol. All distillation experiments use LoRA Hu et al. (2022) with rank $128$, $\alpha=128$, and dropout probability 0. Our optimization protocol employs a learning rate of $0.0005$, weight decay coefficient of $0.1$, and gradient clipping at norm $1.0$. Training follows a cosine learning rate schedule with warm-up over the first 10% of training, batch size $32$, for $4$ epochs. These values are the result of a systematic hyperparameter sweep using the MATH dataset to find configurations that maximize student performance gain.
Figure 4: For both MMLU and MATH data, we show that antidistillation sampling can bring student accuracies down with relatively little cost to the teacher.
4.2 Controlling the utility-distillability trade-off
The bar plots to the right in Figure˜1 show that antidistillation sampling effectively satisfies the desiderata outlined in §3.2. Specifically, for a fixed reduction in teacher performance, students distilled from sampled traces exhibit substantially lower accuracy compared to those distilled from temperature-scaled traces. These initial results, while compelling, represent just one point in the configuration space as they reflect a particular choice of $\lambda$.
In Figures˜3 and 4, we vary $\lambda$ to characterize the degree of control antidistillation sampling provides over the utility-distillability trade-off. Since our experiments use architecturally distinct student and proxy student models, these results confirm that antidistillation sampling generalizes effectively across model families—a critical property for practical deployment.
Recognizing that model owners typically have limited tolerance for sacrificing utility, we explicitly focus on the high-teacher-performance regime in Figure˜9. Even in this setting, where teacher performance degradation is conservative, we observe meaningful degradation in student accuracy, underscoring the method’s practical efficacy. For example, going from 90% to 89% teacher accuracy leads to the poisoned student dropping from 65% to 56% accuracy, while temperature sampling doesn’t degrade the student’s performance at all.
4.3 Diverse configurations for antidistillation sampling
To probe the efficacy of antidistillation sampling across diverse scenarios, we conduct experiments with various teacher-student configurations and datasets. Our primary setup uses Qwen teacher and proxy student models while using a Llama model for the student. We also investigate settings where all the models (student, proxy, and teacher) belong to the same architecture family—either all from the Llama architecture family or all from the Qwen architecture family—evaluated on GSM8K. Results are provided in Figure 10.
Beyond architectural variations, we also validate our finite difference approximation by comparing it with the theoretically-motivated Jacobian-vector product (JVP) implementation (not to be confused with a vector-Jacobian product used in backpropagation). However, even in modern automatic differentiation frameworks, JVPs tend to lack support for a handful of important operations, such as SDPA. Meanwhile, JVPs are more accurately computed in Float32, but flash attention (Dao et al., 2022) only supports Float16 and BFloat16. In our implementation, we then abandon the usage of flash attention, which causes the sampling speed to decrease by around eight times (due to the memory limit, the batch size has to be decreased). Thus, for the rest of our experiments, we use finite differences for the improved convenience and efficiency (See §C for more details). As shown in Figure˜8, both approaches yield similar results in practice, confirming that our computationally efficient finite difference method provides a reliable approximation to the formal gradient-based objective. Further, with our permutation sampling baseline, we provide evidence that $\Delta$ contains useful gradient information by applying random noise with the same statistical properties as $\Delta$ (see §A).
Figure 5: Distillation loss curves show that although the student’s training loss decreases across steps, antidistillation sampling effectively poisons traces, as shown by the increase in the student’s holdout loss.
4.4 Generalizing from the proxy model
Our method demonstrates strong results across various settings, but an important question remains about its underlying mechanism. Since we sample tokens explicitly designed to be detrimental for the proxy model on our holdout set, we are relying on generalization to an unknown student model. Figure˜5 provides insight into this mechanism by tracking loss dynamics during the distillation process. We observe exactly the intended effect. Distillation on the antidistillation traces lowers the student’s loss on the training set while increasing its loss on the holdout set. This result confirms that antidistillation sampling creates traces that are learnable but poison the student model’s ability to reason on the downstream task.
@@ 5 Conclusion
Broader impact.
We expect that antidistillation sampling will have a positive impact on the security of frontier models. By providing a mechanism to protect against distillation, we hope to encourage the continued development of frontier large language models and their applications.
@@ Appendix A Additional Baselines
Appendix B Hyperparameters
B.1 Verifying finite difference approximation
We empirically verify that the finite difference in eq.˜10 behaves as expected by computing the relative error between the finite difference result and term produced from autograd. As shown in fig.˜7, we see it well approximates the autograd computed result for appropriately chosen step size. Here we compute $\left\langle\nabla\ell\left(\theta_{P}\right),\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})\right\rangle$ and stack the different values of $x_{t+1}$ into a $V$ dimensional vector $\widehat{\Delta}$ and compare to the autograd vector $\Delta$. We compute relative error being sensitive only to the direction as
@@ B.1 Verifying finite difference approximation
We run this numerical experiment using Qwen/Qwen2.5-3B. Here we demonstrate that the finite difference can be used to estimate the derivatives in the low precision bfloat16 format. In particular, too small an $\epsilon$ leads to round-off error in the perturbation and too large $\epsilon$ leads to high truncation error in the Taylor expansion, with a sweet spot in the middle. The actual choice of $\epsilon$ may depends heavily on the model size (and numerical precision), so we recommend choosing this value on the exact model in question. In our actual experiment, we pick $\epsilon$ empirically to be $10^{-2}$ as suggested by fig.˜7.
@@
B.2 How We Made The Graphs
[…]
Appendix C Jacobian-Vector Product and Finite Difference
Let us recall that in our derivation, we start from the desired objective function and lead to a Jacobian-vector product (JVP) form:
$\lim_{\eta\to 0}\frac{1}{\eta}\Delta(x_{t+1}|x_{1:t})$ $=\left\langle\nabla\ell\left(\theta_{P}\right),\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})\right\rangle.$ (12)
We then show that due to the symmetry of inner product, this JVP can be approximated by the following finite difference method:
$\widehat{\Delta}(\>\cdot\>|x_{1:t})$ $=\frac{\log p(\>\cdot\>|x_{1:t};\theta_{P}+\epsilon\nabla\ell(\theta_{P}))-\log p(\>\cdot\>|x_{1:t};\theta_{P}-\epsilon\nabla\ell(\theta_{P}))}{2\epsilon}.$ (13)
There are many practical considerations that prevent us from efficiently implementing a memory-friendly JVP computations—for instance, the lack of support for popular attention kernels. At the same time, JVP is better computed in Float32 precisions.
Figure 8: JVP vs. finite difference approximation in antidistillation sampling, evaluated on GSM8k.
In Figure˜8, we further show that even using Float32 precision, finite difference approximation can still outperform JVP in AD sampling.
Figure 9: We zoom in to the first 5% delta of teacher accuracy; these results may be the most practical.
Figure 10: Antidistillation sampling works for a variety of choices of teacher and student models; in the main text, we present results for Qwen teacher and Llama student models. Here, we use either Qwen (left) or Llama (right) for both the teacher and the students; the results remain broadly similar.
Appendix D Example Traces